OpenStudy (anonymous):

A 68 A 68-g bar of gold is cut into 3 equal pieces how does the density of each piece compare to the density of the original gold bar? please i really need help

OpenStudy (anonymous):

Density is mass per volume. each piece now has 1/3 the mass and 1/3 the volume.

OpenStudy (anonymous):

There is no difference. For example, the density of water is 1g/cm^3, whether it is a cup of water or an ocean of water. If the 90g bar of gold of volume 30cm^3. The density will be 90/30=3g/cm^3for example was cut into three pieces, it also means that the mass of each piece would now be 90/3=.30g Also the volume of each piece will be 30/3=10cm^3. From the calculation, the density of each piece will now be 30g/10cm^3, and that is equal to 3g/cm^3, JUST LIKE THE ORIGINAL BAR!!!

OpenStudy (shane_b):

In short the density will not change for the reasons @fall2012 stated. You can prove this pretty easily..start with: \[density=\frac{mass}{volume}\]Since the gold bar is homogenous it has the same density throughout it so dividing its volume into 3 pieces means you're also dividing its mass by 3: \[density=\frac{\frac{mass}{3}}{\frac{volume}{3}}\]\[density=\frac{mass}{\cancel {3}}*\frac{\cancel{3}}{volume}=\frac{mass}{volume}\]